kw= 1.00x10^-14
Determine the equilibrium constant for the following reaction: 4 H,0 = 4H+ + 4 OH
Answer #1

the above reaction is....

4H2O \rightleftharpoons 4H+ + 4OH-

From the above reaction we can say [H+]=[OH-], let say x

So for the above reaction equilibrium constant = \frac{[H+]4*[OH-]4}{[H2O]4}

But the [H2O] will be neglected because its a pure liquid.

So now equilibrium constant will be = [H+]4*[OH-]4

=X4 * X4

=X8

also Kw = [H+]*[OH-] = 1.00 *10-14

OR, X * X =1.00*10-14

X2 = 1.00*10-14

Equilibrium constant = X8

= (X2)4

=(1.00*10-14)4

=1.00*10-56